OpenStudy (anonymous):

equation of a circle S(x,y)=0 , (S(2,3)=16) which touches the line 3x+4y-7=0 at (1,1) is given by A) x^2+y^2+2X+2Y-6=o B)x^2+y^2+X+Y=0

OpenStudy (anonymous):

solve the linear equation for x or y and sub into the circle equation. i.e x=(7-4y)/3 replace that as x in the circle and then it can be solved

OpenStudy (anonymous):

What does (2,3)=16 mean?

OpenStudy (ankit042):

when you put the point (2,3) in the equation of circle you get 16

OpenStudy (anonymous):

Thank you ankit and taplin

OpenStudy (ankit042):

@Ankitha007 have yous solved the question. I am sure how to proceed with Taplin's method

OpenStudy (anonymous):

I am solving ... :) I've understood its A) answer I hope.

OpenStudy (ankit042):

i will assume the eq of circle be (x-a)^2+(y-b)^2=r^2 (a,b,r) 3 unknowns now we can make 3 equations f(2,3)=16 f(1,1) =0 and b-1/a-1=4/3

OpenStudy (ankit042):

can someone help me understand the method used by @Taplin44

OpenStudy (anonymous):

if the line and circle intersect then the x and y values must be the same.

OpenStudy (anonymous):

at that point anyway

OpenStudy (ankit042):

@Taplin44 Yes at the point of contact values will be same... but not at all the points! I am not sure you can replace value from linear equation

OpenStudy (ankit042):

And question says it touches at (1,1) not intersect so the line will be a tangent

OpenStudy (anonymous):

Any which touches 3X+4y-7=o will be of the form S(x,y) = (x-1)^2 +(y-1)^2 + I ( 3X+4Y-7) = 0 u said S(x,y)=16 therefore I=1 therefore we get circle equation A)

OpenStudy (anonymous):

@ankit042 uve understood ....

OpenStudy (ankit042):

@Ankitha007 I guess I can recall this formula now...just thinking of the proof. The method I mentioned will work but this one will be faster! Thanks

OpenStudy (anonymous):

Yes just studied this but dint pass through my mind ... Thanx

OpenStudy (dan815):

1,1 is a point on the circle only equation a satisfies this condition xD

OpenStudy (dan815):

ya probably best to get familiar with this method though.. if you want to actually learn "i will assume the eq of circle be (x-a)^2+(y-b)^2=r^2 (a,b,r) 3 unknowns now we can make 3 equations f(2,3)=16 f(1,1) =0 and b-1/a-1=4/3"

OpenStudy (anonymous):

Yeah to understand yes